OpenStudy (anonymous):

What is the period of f(x) = - tan 4x?

OpenStudy (ash2326):

@Pssssst Do you know the period of tan x?

OpenStudy (anonymous):

No, I do not.

OpenStudy (ash2326):

Ok, do you know what does period mean ?

OpenStudy (ash2326):

For a function

OpenStudy (anonymous):

Honestly, no.

OpenStudy (ash2326):

OK, I'll explain :)

OpenStudy (anonymous):

Thank you, and sorry haha. :)

OpenStudy (ash2326):

Let a function be f(x), T is it's period if \[f(x+T)=f(x)\] it'll also hold for integral multiples of T \[f(x+nT)=f(x)\] n=1, 2, 3,4, 0, -1, -2

OpenStudy (ash2326):

If you look at the graph of tan x, it's period is \(2\pi\)

OpenStudy (ash2326):

Do you get this?

OpenStudy (anonymous):

Alright. I'm following.

OpenStudy (ash2326):

Now if we have tan 2x, it's period will be \(\frac{2\pi}{2}\) can you tell me, what will be the period of tan 4x ?

OpenStudy (ash2326):

@Pssssst ??

OpenStudy (anonymous):

I'm sorry, I had to do something. hah

OpenStudy (anonymous):

4x/2

OpenStudy (anonymous):

My bad, * 4x/4 right?

OpenStudy (ash2326):

Read carefully at my posts again :)

OpenStudy (anonymous):

The period of tan 2x is 2(pi)/2? or 2x/2? It's hard to read it on my computer